### v3

@@ 4.3 Comparison to existing methods
In this section we compare the performance of LaDynS on simulated data generated from the shared oscillatory driver model (see Section 4.2) with three existing methods: Dynamic Kernel Canonical Correlation Analysis (DKCCA, Rodu et al., 2018), Delayed Latents Across Groups (DLAG, Gokcen et al., 2022), and Gaussian Process Factor Analysis (GPFA, Yu et al., 2009). The simulated data are generated under two distinct connectivity scenarios:
@@
All other tuning parameters followed the example codes provided by each software package.
@@
Under the single-factor scenario, where the interaction between a single pair of latent factors is sufficiently strong, GPFA was able to recover all three interaction epochs (Figure 11(c-e)), matching the performance of LaDynS (Figure 11(a), center panel). By contrast, DLAG struggled to capture the bidirectional nature of the interaction (Figure 11(b)), consistent with its assumption that each pair of between-group latent factors has a fixed lead-lag relationship. DKCCA again did not detect any meaningful interaction in this setting.
@@ 4.4 Experimental Data Analysis
We applied LaDynS to local field potentials (LFPs), collected in the experiment described in Khanna et al. (2020), from two Utah arrays implanted in a Macaque monkey’s prefrontal cortex (PFC) and V4 during a memory-guided saccade task. Each trial of the task started with a monkey fixating its eyes on a dot at the center of the screen. A visual cue was given for $50$ ms to indicate a target location, which was randomly chosen from $40$ candidate locations (8 directions and 5 amplitudes) that tiled the display screen. The cue was turned off and the monkey had to remember the target location while maintaining eye fixation for a delay period of $500$ ms. After the delay period, the monkey made a saccade to the remembered position, and a liquid reward was provided on successful trials. As in Khanna et al. (2020), we analyzed the time series during the delay period, based on $3000$ successful trials. Time $t=0$ corresponds to the start of that period. The data are available in Snyder et al. (2022). Because beta oscillations are often associated with communication across brain areas (Klein et al., 2020; Miller et al., 2018), we filtered LFP recordings at a beta oscillation frequency $18$ Hz and obtained the beta oscillation power envelopes as described in Section 4.2.1. We chose $18$ Hz because it was the frequency having the largest power within the range 12-40 Hz (see Figure S11). After downsampling the power envelopes to $200$ Hz, we applied LaDynS with the regularizer $\lambda_{\text{diag}}$ on the diagonal of $\Sigma$, as in Section 4.2.1, because the filtered data were very smooth.
@@ 5 Discussion
The partial $R^{2}$ values in Figure 12(b) are highly statistically significant, but they are very small. We note that our diagonal regularization of the estimated latent precision matrix artificially reduced these values, so that the scale is no longer interpretable in familiar terms. The cross-correlation estimates are several orders of magnitude larger (see Figure S12). However, they remain smaller than .1, as are the raw correlations at 400 ms and 425 ms, respectively, across individual electrodes in V4 and PFC (not shown). This reflects the highly inhomogeneous nature of the LFP voltage recordings together with the dominance of low frequencies (unfiltered LFP spectra typically follow $1/f^{k}$ trends, where $f$ is frequency and $k$ is roughly 1.5 to 2; see Kass et al., 2023). The large number of repeated trials allowed us to find and document the results.
@@
In addition to making the analysis possible, the repeated trial structure suggests substantive interpretation based on trial-to-trial variability. Although investigators take pains to make the experimental setting nearly the same on each trial, the inevitable small fluctuations in the way the subject interacts with the environment, together with changes in the subject’s underlying state (involving alterations in motivational drive, for example), lead to observable fluctuations in behavior and in the recorded neural activity. Although the network sources of trial-to-trial variability in the PFC and V4 data are unknown, they produce the kind of correlated activity revealed in Figure 12(b). To interpret it, we acknowledge there could be some trial-to-trial variation (such as changing motivational state) in the inputs that drive beta power in V4 and PFC, having just the right differential time lags to produce the correlated activity picked up by the partial $R^{2}$ values. Could such task-irrelevant pulses of activity change across time, within repetitions of the task, in such a way as to produce, the peaks in Figure 12(b)? It is possible, but it would be surprising, especially when we consider contemporary ideas about beta oscillations during working memory tasks (Miller et al., 2018) along with the well-identified distinction between early and late visual processing that, presumably, corresponds to a distinction between feedforward and recurrent (bidirectional) flow of information (Chen et al., 2022; Del Cul et al., 2007; Supèr et al., 2001; Yang et al., 2019). The alternative we mentioned, that PFC and V4 are involved, together, in goal-directed visual processing and memory, with the two areas acting bidirectionally around 400 ms post stimulus, seems likely.
@@
There are many ways to extend the ideas developed here. For band-pass filtered data, such as those analyzed in Section 4.4, phase analysis (Klein et al., 2020) could be combined with amplitude analysis based on the complex normal distribution, as in Urban et al. (2023) (which developed a latent variable model at a single time point). Multiple frequencies could be considered, along the lines of Gallagher et al. (2017), as well. In Bong et al. (2020) we have described an extension of LaDynS, where the within-region noise vectors $\epsilon_{k}$ were allowed to have general time series structure and the latent time series driving each brain region were multidimensional. That brief report (using data filtered at a different frequency) does not provide the details given here, nor does it include inferential procedures. A future step would be to generalize the methods here to the setting of Bong et al. (2020). A different direction for additional research would be to simplify the version of LaDynS we have used here by imposing suitable spatiotemporal structure on the latent time series. Regardless of whether such approaches are fruitful, the general framework of LaDynS could be of use whenever interest focuses on time-varying interactions among groups of repeatedly-observed multivariate neural time series.
@@
which is maximized when $\Omega^{\prime}=\overline{\Sigma}^{{}^{\prime}-1}$ given $w_{k}^{\prime(t)}$’s are fixed. sThus, the maximum likelihood estimation is equivalent to finding $w_{k}^{\prime(t)}$ minimizing $\log\det(\overline{\Sigma}^{\prime})$ under $w_{k}^{\prime(t)\top}\overline{S}_{kk}^{(t,t)}w_{k}^{\prime(t)}=1$ for $k\in[K]$, which is the GENVAR procedure of Kettenring (1971). This proves the desired results for the MLE with $m_{k}^{(t)}=1$. The other MLEs with $m_{k}^{(t)}<1$ correspond to the non-identifiable parameter sets, which have $u_{k}^{(t)\top}u_{k}^{(t)}<1$.
@@
To update $\Omega$, we use the P-gLASSO algorithm of Mazumder and Hastie (2012), which is more efficient than the original gLASSO algorithm of Friedman et al. (2008). The efficiency is attributed to P-gLASSO’s flexibility with initial values, whereas gLASSO operates with a strict choice of initial $\widehat{\Omega}$ ($\overline{\Sigma}^{-1}$ in case of Eq. 11). In Algorithm S1, the estimate $\widehat{\Omega}$ from the past iteration serves as a warm start for the next iteration, so that we do not have to redo the entire paths from $\overline{\Sigma}^{-1}$ to $\widehat{\Omega}$. We can further reduce the computation cost by harnessing the banded sparse structure of $\Omega$ in Fig. 4. In Algorithm S2, we provide a modified P-gLASSO algorithm designed for the banded sparsity. The modified algorithm reduces the size of the LASSO sub-problem from $2T$ to $2d_{\text{cross}}+2d_{\text{auto}}$ and the computational cost of a P-gLASSO iteration from $O(T^{4}+T^{3}N)$ to $O(T((d_{\text{cross}}+d_{\text{auto}})^{3}+(d_{\text{cross}}+d_{\text{auto}})^{2}N)$.

### v4

@@ 4.3 Comparison to existing methods
In this section we compare the performance of LaDynS on simulated data generated from the shared oscillatory driver model (see Section 4.2) with three existing methods: Dynamic Kernel Canonical Correlation Analysis (DKCCA, rodu2018detecting), Delayed Latents Across Groups (DLAG, gokcen2021disentangling), and Gaussian Process Factor Analysis (GPFA, yu2009gaussian). The simulated data are generated under two distinct connectivity scenarios:
@@
For all other tuning hyperparameters of DKCCA, DLAG, and GPFA, we used the default settings provided in the example scripts of the respective code packages. For details, we refer the reader to the example code in our ladyns package.
@@
Under the single-factor scenario, where the interaction between a single pair of latent factors is sufficiently strong, GPFA was able to recover all three interaction epochs (Figure 11(c-e)), matching the performance of LaDynS (Figure 11(a), center panel). By contrast, DLAG struggled to capture the bidirectional nature of the interaction (Figure 11(b)), consistent with its assumption that each pair of between-group latent factors has a fixed lead-lag relationship. We additionally ran DLAG with a single between-group latent factor to match the true latent dimensionality in this scenario; however, its qualitative behavior remained unchanged, and DLAG continued to struggle to capture the bidirectional nature of the interaction. DKCCA again did not detect any meaningful interaction in this setting.
@@ 4.4 Experimental Data Analysis
We applied LaDynS to local field potentials (LFPs), collected in the experiment described in khanna2019dynamic, from two Utah arrays implanted in a Macaque monkey’s prefrontal cortex (PFC) and V4 during a memory-guided saccade task. Each trial of the task started with a monkey fixating its eyes on a dot at the center of the screen. A visual cue was given for $50$ ms to indicate a target location, which was randomly chosen from $40$ candidate locations (8 directions and 5 amplitudes) that tiled the display screen. The cue was turned off and the monkey had to remember the target location while maintaining eye fixation for a delay period of $500$ ms. After the delay period, the monkey made a saccade to the remembered position, and a liquid reward was provided on successful trials. As in khanna2019dynamic, we analyzed the time series during the delay period, based on $3000$ successful trials. Time $t=0$ corresponds to the start of that period. The data are available in snyder2022utah. Because beta oscillations are often associated with communication across brain areas (klein2020torus; miller2018wm2), we filtered LFP recordings at a beta oscillation frequency $18$ Hz and obtained the beta oscillation power envelopes as described in Section 4.2.1. We chose $18$ Hz because it was the frequency having the largest power within the range 12-40 Hz (see Figure S11). After downsampling the power envelopes to $200$ Hz, we applied LaDynS with the regularizer $\lambda_{\text{diag}}$ on the diagonal of $\Sigma$, as in Section 4.2.1, because the filtered data were very smooth.
@@ 5 Discussion
The partial $R^{2}$ values in Figure 12(b) are highly statistically significant, but they are very small. We note that our diagonal regularization of the estimated latent precision matrix artificially reduced these values, so that the scale is no longer interpretable in familiar terms. The cross-correlation estimates are several orders of magnitude larger (see Figure S12). However, they remain smaller than .1, as are the raw correlations at 400 ms and 425 ms, respectively, across individual electrodes in V4 and PFC (not shown). This reflects the highly inhomogeneous nature of the LFP voltage recordings together with the dominance of low frequencies (unfiltered LFP spectra typically follow $1/f^{\alpha}$ trends, where $f$ is frequency and $\alpha$ is roughly 1.5 to 2; see kass2023identification). The large number of repeated trials allowed us to find and document the results.
@@
In addition to making the analysis possible, the repeated trial structure suggests substantive interpretation based on trial-to-trial variability. Although investigators take pains to make the experimental setting nearly the same on each trial, the inevitable small fluctuations in the way the subject interacts with the environment, together with changes in the subject’s underlying state (involving alterations in motivational drive, for example), lead to observable fluctuations in behavior and in the recorded neural activity. Although the network sources of trial-to-trial variability in the PFC and V4 data are unknown, they produce the kind of correlated activity revealed in Figure 12(b). To interpret it, we acknowledge there could be some potentially confounding trial-to-trial variation — such as changing motivational state, stimulus timing, bandpass-filter distortions and volume conduction — that drive beta power in V4 and PFC, having just the right differential time lags to produce the correlated activity picked up by the partial $R^{2}$ values, as cautioned by lintas2021operant. Could such task-irrelevant pulses of activity change across time, within repetitions of the task, in such a way as to produce, the peaks in Figure 12(b)? It is possible, but it would be surprising, especially when we consider contemporary ideas about beta oscillations during working memory tasks (miller2018wm2) along with the well-identified distinction between early and late visual processing that, presumably, corresponds to a distinction between feedforward and recurrent (bidirectional) flow of information (super2001two; buschman2007topdown; del2007brain; yang2019exploring; chen2022population; mejias2016feedforward). The alternative we mentioned, that PFC and V4 are involved, together, in goal-directed visual processing and memory, with the two areas acting bidirectionally around 400 ms post stimulus, seems likely.
There are many ways to extend the ideas developed here. While we applied LaDynS to LFP recordings, the framework is applicable more broadly to other slowly varying multidimensional time series. In particular, the proposed methodology can be extended to other neural modalities such as EEG, MEG, and fMRI. Although data from these modalities often violate global stationarity assumptions underlying classical approaches such as cross-correlograms (guan2020profiles; klonowski2009everything; chang2010time), local stationarity remains reasonable when functional connectivity depends on cognitive/behavioral states which evolve on time scales longer than the sampling interval. While we have not yet explored such extensions empirically, LaDynS is particularly appealing in this regime, as it explicitly accommodates slowly varying connectivity and cross-correlation over time. At the same time, signals in EEG and MEG often reflect mixtures of activity from multiple brain structures and switches between metastable neural states, which may introduce additional confounding and complicate the interpretation of inferred associations.
LaDynS may perform less favorably in settings where within-region noise autocorrelations dominate the signal. The simulation results in Figure 4 of bong2020latent illustrate the resulting loss of specificity under such adversarial conditions. In these cases, methods that explicitly model within region autocorrelated noise, such as DLAG (gokcen2021disentangling), can outperform LaDynS. In bong2020latent, we proposed an extension of LaDynS that addresses this issue by allowing the within region noise processes $\epsilon_{k}$ to follow general time series structures and by modeling the latent processes driving each brain region as multidimensional. That brief report, which focused on data filtered at a different frequency band, did not provide the methodological details or inferential procedures developed here. An important direction for future work is to generalize the present framework to the setting considered in bong2020latent.
Another important direction for future research is a theoretical analysis of how regularization hyperparameters affect recovery of the underlying lag structure. In this work, we focused on proposing a regularization based approach to functional connectivity estimation together with a data driven hyperparameter calibration scheme and demonstrated its performance through simulation studies, while establishing theoretical guarantees for hyperparameter selection and optimality remains an interesting topic for future investigation. In addition, developing theory that accounts for dependence across experimental trials is of particular interest, as such dependence may arise from long term memory or adaptation effects in neural recordings and could impact the validity of the proposed inference procedure.
@@
Furthermore, for band-pass filtered data, such as those analyzed in Section 4.4, phase analysis (klein2020torus) could be combined with amplitude analysis based on the complex normal distribution, as in urban2023oscillating (which developed a latent variable model at a single time point). Multiple frequencies and nonlinear association, such as cross-bicoherence, could be considered, along the lines of tort2010measuring; gallagher2017cross; aliramezani2024delta; abe2024detection, as well. A different direction for additional research would be to simplify the version of LaDynS we have used here by imposing suitable spatiotemporal structure on the latent time series. Regardless of whether such approaches are fruitful, the general framework of LaDynS could be of use whenever interest focuses on time-varying interactions among groups of repeatedly-observed multivariate neural time series.
@@
which is maximized when $\Omega^{\prime}=\overline{\Sigma}^{{}^{\prime}-1}$ given $w_{k}^{\prime(t)}$’s are fixed. sThus, the maximum likelihood estimation is equivalent to finding $w_{k}^{\prime(t)}$ minimizing $\log\det(\overline{\Sigma}^{\prime})$ under $w_{k}^{\prime(t)\top}\overline{S}_{kk}^{(t,t)}w_{k}^{\prime(t)}=1$ for $k\in[K]$, which is the GENVAR procedure of kettenring1971canonical. This proves the desired results for the MLE with $m_{k}^{(t)}=1$. The other MLEs with $m_{k}^{(t)}<1$ correspond to the non-identifiable parameter sets, which have $u_{k}^{(t)\top}u_{k}^{(t)}<1$.
@@
To update $\Omega$, we use the P-gLASSO algorithm of mazumder2012graphical, which is more efficient than the original gLASSO algorithm of friedman2008sparse. The efficiency is attributed to P-gLASSO’s flexibility with initial values, whereas gLASSO operates with a strict choice of initial $\widehat{\Omega}$ ($\overline{\Sigma}^{-1}$ in case of Eq. 11). In Algorithm S1, the estimate $\widehat{\Omega}$ from the past iteration serves as a warm start for the next iteration, so that we do not have to redo the entire paths from $\overline{\Sigma}^{-1}$ to $\widehat{\Omega}$. We can further reduce the computation cost by harnessing the banded sparse structure of $\Omega$ in Fig. 4. In Algorithm S2, we provide a modified P-gLASSO algorithm designed for the banded sparsity. The modified algorithm reduces the size of the LASSO sub-problem from $2T$ to $2d_{\text{cross}}+2d_{\text{auto}}$ and the computational cost of a P-gLASSO iteration from $O(T^{4}+T^{3}N)$ to $O(T((d_{\text{cross}}+d_{\text{auto}})^{3}+(d_{\text{cross}}+d_{\text{auto}})^{2}N)$.